OpenStudy (anonymous):

Combine the terms. 7/8x-1/2x A.1/8x B. 3/8 C. 3/8x D. x @iGreen @Catlover5925 @Nnesha @monsterfox1997 @chefducky @fallenangelorchid please help fan and medal

OpenStudy (p0sitr0n):

1/2=4/8 7/8x-4/8x=3/8x